OpenStudy (shubhamsrg):

if (b^4) + (1/b)^4 =6 then find the value of (b+ i/b)^16 where i = sqrt(-1)

OpenStudy (shubhamsrg):

@mukushla ??

OpenStudy (anonymous):

use this \[(b^2-\frac{1}{b^2})^2+2=b^4+\frac{1}{b^4}=6\]\[(b^2-\frac{1}{b^2})^2=4\]\[b^2-\frac{1}{b^2}=(b+\frac{i}{b})^2-2i=\pm2\]

OpenStudy (shubhamsrg):

omg!! buddy,,which planet are you from? seriously??

OpenStudy (shubhamsrg):

thanks a lot!

OpenStudy (anonymous):

lol....yw :)